OpenStudy (anonymous):

find the circumfrence of the circle pictured. Use C=3.14(2)(r).

OpenStudy (anonymous):



Parth (parthkohli):

2x + 1 is the radius right?

OpenStudy (anonymous):

a. 6.28+3.14 b.^2+6.28 c.12.56x+6.28 and yes Parth Kohli:)

Parth (parthkohli):

2 * (2x + 1) * 3.14 = 6.28(2x + 1) = 12.56x + 6.28 -> Option C

OpenStudy (anonymous):

thanks for all of your help!

Parth (parthkohli):

:)